OpenStudy (anonymous):

x-y=18 3/4 x + 3/4 y = -15/2 how do i solve this linear system with subsitution ?

OpenStudy (anonymous):

3/4 x + 3/4y = -15/2 multiplied both sides by 4/3 >> x + y = -60/6 = -10 then the rest should be easy

OpenStudy (anonymous):

i have no clue how you got there . where did the -10 come from? shouldnt it be multiplied by 3/4?

OpenStudy (saifoo.khan):

still stuck?

OpenStudy (anonymous):

-60 over 6 equals -10 Then you got two equations x-y=18 x + y = -10

OpenStudy (anonymous):

yes im still stuck .

OpenStudy (anonymous):

x - y = 18 x + y = -10  2x = 8 x = 4

OpenStudy (saifoo.khan):